​Allied Industries, Inc. Selected Income Statement Items, 2017 Cash Sales​$1,500,000 Credit Sales​$7,500,000 Total Sales​$9,000,000 COGS​$6,000,000 ​Allied Industries, Inc. ​Selected Balance Sheet Accounts ​12/31/2017​12/31/2016​Change Accounts Receivable​$270,000​$240,000​$30,000 Inventory​$125,000​$100,000​$25,000 Accounts Payable​$110,000​$90,000​​​$20,000 45) Using the information provided, what is the length of the production cycle for the firm? A) 6.08 days B) 7.60 days C) 53.33 days D) 6.84 days 46) Using the information provided, what is the accounts receivable turnover for the firm? A) 23.53 times B) 29.41 times C) 53.33 times D) 60.00 times 47) Using the information provided, what is the collection cycle for the firm? A) 6.84 days B) 7.60 days C) 10.34 days D) 12.41 days 49) Using the information provided, what is the accounts payable cycle for the firm? A) 4.06 days B) 4.87 days C) 6.08 days D) 24.33 days
Answer #1

Answer

45) Option 'D '-6.84 days is the correct option

working note

Length of production cycle = (365*Average invetory)/ Cost of Goods Sold
Average inventory =(125,000+100,000)/2
=112500
Cost of Goods Sold =$6,000,000
So ,length of production cycle =(365*112500)/6,000,000
=6.84 days

46) Option 'B ' -29.41 times is the correct option

working note

Accounts recevable turnover = Credit sales / Average Accounts Receivable
Credit sale =$7,500,000
Average Acccounts Receivable =($270,000+$240,000) /2
=255,000
Accounts recevable turnover =7,500,000 /255,000
   =29.41

47) Option 'D' -12.41 days' is the correct option

working note

Collection cycle = 365 /Accounts receivable turnover
=365 / 29.41
=12.41 days

49)Option 'C '-6.08 days is the correct option

working note

Accounts Payable cycle =(Average Accounts payable / Cost of Goods Sold) *365
Average Accounts payable =(110,000+90,000) /2
=100,000
Cost of Goods Sold =$6,000,000
So ,Accounts Payable cycle =(100,000 /6,000,000)*365
6.08